40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL, Page 23 of 52  Not logged in ELOG logo
ID Date Author Type Category Subject
  1501   Thu Sep 4 11:47:42 2014 EvanDailyProgressNoiseBudgetAlGaAs python noise budget: photothermal TFs added

The first attachment shows the photothermal TFs which take absorbed power (in watts) to the mirror displacement (in meters) as sensed by our 215-µm beam. Since last night, I've fixed the coating TE part and committed the updated ipynb to the SVN.

The second attachment shows the noise budget, with the photothermal shot noise contribution.

Attachment 1: photothermalTF.pdf
photothermalTF.pdf
Attachment 2: noiseBudget.pdf
noiseBudget.pdf
  1500   Wed Sep 3 16:34:37 2014 Tara, EvanDailyProgressBEATBeat, mode-matching

Quote:

Tara added some more juice to the north cavity heater last night. Now we can lock both cavities to TEM00 and get a beat within the bandwidth of the 1811.

  • North laser slow: 5.020 V
  • South laser slow: 0.722 V
  • Beat frequency: 49.3 MHz

Beat frequency drifted to 61 MHz over the course of a few hours. We need to wait for the cavity temperatures to settle.

I improved the mode-matching a little bit on the south cavity; it's about 50% (the theoretical max is 71%). The south lenses are now on translation stages.

I've attached a beat spectrum. Nothing is floated, RAM is not optimized, etc.; this is just a rough indicator of where things stand.

Here is what I think should happen next, in rough order of importance:

  1. Float chamber
  2. Measure RIN
  3. Measure photothermal TF (I also need to recheck my photothermal code — I don't believe the coating TE part)
  4. Put photothermal noise on noise budget
  5. Reduce RAM.
  6. Measure residual frequency noise.
  7. Measure PLL noise. Use ATF DAQ and make spectral histogram.
  8. Measure seismic noise (with Guralp or T240), with table floated and unfloated. Use ATF DAQ and make spectral histogram.
Attachment 1: noiseBudget.pdf
noiseBudget.pdf
  1499   Wed Sep 3 12:02:19 2014 Tara, EvanDailyProgressBEATBeat found

Tara added some more juice to the north cavity heater last night. Now we can lock both cavities to TEM00 and get a beat within the bandwidth of the 1811.

  • North laser slow: 5.020 V
  • South laser slow: 0.722 V
  • Beat frequency: 49.3 MHz
  1498   Tue Sep 2 17:11:27 2014 EvanDailyProgressBEATNo beat

Quote:

Quote:

Searched around over various axial modes in order to find a beat.

I fiddled a bit with the output QWPs in order to get the polarizations to match. Because of the birefringent coatings, light transmitted through the cavity is not circular, and the polarization state will depend on which of the two modes we lock to. In case, the original QWP angles were 202° for north and 19° for south.

On Koji's suggestion, I set up a second 1811 to monitor the beat on the input side of the cavities, so that we can see what the lasers are doing independent of the cavity resonances. For each path, I am using the s-polarized light that is rejected from a PBS, so that we don't need to add extra optics to the beam paths.

For example, for south slow at 1.206 V and north slow at 5.565 V, I get a 69 MHz beat (with both cavities unlocked).

We should use this in conjunction with the cavity locking optics to figure out what the correct axial cavity modes are, and whether the cavities need any temperature adjustment.

With this setup, I find that a sub-100-MHz beat occurs for north slow at 5.015 V and south slow at 0.725 V. This south slow voltage corresponds to a south cavity TEM00 mode, but the nearest north slow voltage is at 5.298 V.

  1497   Tue Sep 2 15:04:05 2014 EvanDailyProgressBEATNo beat

Quote:

Searched around over various axial modes in order to find a beat.

I fiddled a bit with the output QWPs in order to get the polarizations to match. Because of the birefringent coatings, light transmitted through the cavity is not circular, and the polarization state will depend on which of the two modes we lock to. In case, the original QWP angles were 202° for north and 19° for south.

On Koji's suggestion, I set up a second 1811 to monitor the beat on the input side of the cavities, so that we can see what the lasers are doing independent of the cavity resonances. For each path, I am using the s-polarized light that is rejected from a PBS, so that we don't need to add extra optics to the beam paths.

For example, for south slow at 1.206 V and north slow at 5.565 V, I get a 69 MHz beat (with both cavities unlocked).

We should use this in conjunction with the cavity locking optics to figure out what the correct axial cavity modes are, and whether the cavities need any temperature adjustment.

  1496   Tue Sep 2 11:06:54 2014 EvanDailyProgressBEATNo beat

Searched around over various axial modes in order to find a beat.

I fiddled a bit with the output QWPs in order to get the polarizations to match. Because of the birefringent coatings, light transmitted through the cavity is not circular, and the polarization state will depend on which of the two modes we lock to. In case, the original QWP angles were 202° for north and 19° for south.

  1495   Sat Aug 30 19:10:11 2014 EvanDailyProgressNoiseBudgetAlGaAs python noise budget: seismic coupling added

Same data and same isolation model as for the silica/tantala noise budget. Since we have new table legs, we should retake this data (and make a spectral histogram).

The resonance frequencies of the stack are given as 10 Hz and 35 Hz in the noise budget. Are these for the old stack? I recall that with the new stack we measured resonances at 3, 7, and 10 Hz.

Also I want to double check the sequence of interpolation steps we've used on the silica/tantala noise budget. There are some seismic peaks and silica/tantala beat peaks that almost (but don't quite) match up in frequency, and I wonder whether this is an artifact of the interpolation.

Attachment 1: noiseBudget.pdf
noiseBudget.pdf
  1494   Fri Aug 29 19:56:06 2014 EvanDailyProgressBEATBeat breadboard in place

Quote:

Beat breadboard is slid back into place. North transmission appears on north camera. Still need to do south transmission.

Tara has found south transmission on camera. I steered the transmitted beams onto the beat PD and then made the k-vectors as parallel as I could as seen on an IR card.

The DC voltage on the PD is okay (ca. 50 mV from each beam), but I cannot see a beat note on the AC path using the HP4395. Tara will give a temperature kick which hopefully will bring the beat note within the range of the 1811.

  1493   Fri Aug 29 15:35:55 2014 EvanDailyProgressBEATMode-matching for beat

I predict (via alm) that the spot size on the diode (z = 991 mm) is 79 µm in the current configuration.

Attachment 1: ctnbeat_algaas.pdf
ctnbeat_algaas.pdf
Attachment 2: ctnbeat_algaas_alm.zip
  1492   Fri Aug 29 12:35:54 2014 EvanDailyProgressBEATBeat breadboard in place

Beat breadboard is slid back into place. North transmission appears on north camera. Still need to do south transmission.

  1491   Fri Aug 29 09:21:26 2014 EvanDailyProgressFSSPDH shot noise estimate

Ignoring for the time being the issue of offsets in the PDH error signal, here's my prediction for the new PDH shot noise level, assuming a visibility of 0.92 × 0.7 = 0.64 and an incident power of 2 mW.

So our beat will be slightly worse around 1 kHz, but we aren't completely hosed by the shot noise. I'd think the true solution here is to find (or buy) two large-aperture Faraday isolators to replace the PBS+QWP setup (according to alm, the spot size in this region is about 1.1 mm). E.g., we might consider a large-aperture ThorLabs isolator.

Attachment 1: noiseBudget.pdf
noiseBudget.pdf
  1490   Thu Aug 28 20:47:24 2014 EvanDailyProgressFSSPDH error signal on north cavity

Quote:

North PZT sweep: 10 Vpp triangle wave, 3 Hz

North slow control voltage: 3.6805 V

Actuation on north broadband EOM removed

Phase tuning needed, mode-matching needed

Find TNC-SMA converters

 Here's what I expect to happen given

  • perfect mode-matching,
  • critical coupling with 150 ppm transmissivity for each mirror,
  • p/s mode splitting of 2.0 MHz, and
  • perfect demod phase.

It seems to match up qualitatively with the measurement. In particular, it does not appear possible to exceed 70% visibility for each mode.

Attachment 1: npdh_sim.pdf
npdh_sim.pdf
Attachment 2: npdh_sim_narrow.pdf
npdh_sim_narrow.pdf
Attachment 3: fpbirefringence.pdf
fpbirefringence.pdf
  1489   Thu Aug 28 19:10:40 2014 EvanDailyProgressopticMode-matching solution for north cavity

Quote:

Current configuration:

  • Target waist: 180 µm, z = 0 mm
  • Lens 1: 140 mm focal length, z = −711 mm (24″ from center of vacuum chamber + 4″ through periscope)
  • Lens 2: 84 mm focal length, z = −991 mm (11″ further behind lens 1)
  • Seed waist = ??

Since we know we were mode-matched fairly well into the 180 µm waist of the silica/tantala cavity (>93% visibility), I asked alm to propagate this waist backward through the lenses in order to find a seed waist. It reports a waist of 161 µm at z = −1373 mm.

I asked alm for a new configuration using the same two lenses. The best configuration (mode overlap = 1) is as follows:

  • Seed waist: 161 µm at z = −1373 mm
  • Lens 1: 140 mm focal length, z = −743 mm
  • Lens 2: 84 mm focal length, z = −1023 mm
  • Target waist: 215 µm, z = 0 mm

So we should move lens 1 back by 32 mm (=1.3″), and move lens 2 back by the same amount.

I moved both lens mounts back by 1″, then adjusted the Vernier knobs and periscope mirrors to try to maximize the visibility as seen on north REFL DC.

The best I am able to do so far is a visibility of v = 1 − 0.57(1) V / 1.74(1) V = 0.672(6).

  1488   Thu Aug 28 17:36:03 2014 EvanDailyProgressopticMode-matching solution for north cavity

Current configuration:

  • Target waist: 180 µm, z = 0 mm
  • Lens 1: 140 mm focal length, z = −711 mm (24″ from center of vacuum chamber + 4″ through periscope)
  • Lens 2: 84 mm focal length, z = −991 mm (11″ further behind lens 1)
  • Seed waist = ??

Since we know we were mode-matched fairly well into the 180 µm waist of the silica/tantala cavity (>93% visibility), I asked alm to propagate this waist backward through the lenses in order to find a seed waist. It reports a waist of 161 µm at z = −1373 mm.

I asked alm for a new configuration using the same two lenses. The best configuration (mode overlap = 1) is as follows:

  • Seed waist: 161 µm at z = −1373 mm
  • Lens 1: 140 mm focal length, z = −743 mm
  • Lens 2: 84 mm focal length, z = −1023 mm
  • Target waist: 215 µm, z = 0 mm

So we should move lens 1 back by 32 mm (=1.3″), and move lens 2 back by the same amount.

Attachment 1: ctn_algaas_alm.pdf
ctn_algaas_alm.pdf
Attachment 2: ctn_algaas.zip
  1487   Thu Aug 28 13:40:05 2014 EvanDailyProgressFSSPDH error signal on north cavity

North PZT sweep: 10 Vpp triangle wave, 3 Hz

North slow control voltage: 3.6805 V

Actuation on north broadband EOM removed

Phase tuning needed, mode-matching needed

Find TNC-SMA converters

Attachment 1: npdh_broad.pdf
npdh_broad.pdf
Attachment 2: npdh_fine.pdf
npdh_fine.pdf
Attachment 3: npdh_sweep.zip
  1486   Wed Aug 27 03:21:53 2014 ranaSummaryopticoptimization for ETM with a-Si/SiO2 coatings

 I filled in more values for a-Si at 120 K into the wiki that Matt Abernathy set up. Then I ran the optimization code for Brownian noise only:

 aSi_120_Layers_60000.pdf

The above plot shows the comparison between the optimized aLIGO coating (silica:tantala at 300K) v. the a-Si coating at 120 K.

 aSi_R_60000.pdf

Then, finally, I compared the TO and Brownian noise of the two designs using the plotTO120.m script:

 aSi_120_TOnoise_60000.pdf

The dashed curves are silica:tantala and the solid lines are a-Si:silica. The Brownian noise improvement is a factor of ~6. A factor of ~1.6 comes from the temperature and the remaining factor of ~3.9 comes from the low loss and the lower number of layers.

I think this is not yet the global optimum, but just what I got with a couple hours of fmincon. On the next iteration, we should make sure that we minimize the sensitvity to coating thickness variations. As it turns out, there was no need to do the thermo optic cancellation since the thermo-elastic is so low and the thermo-refractive is below the Brownian almost at all frequencies.

  1485   Mon Aug 25 20:54:38 2014 EvanNotesNoiseBudgetAlGaAs python noise budget: TO implemented

Quote:

Quote:

I have started a python implementation of the AlGaAs noise budget. All parameters, functions, etc. are defined in a single notebook, and this same notebook generates the plot. The python uncertainties package facilitates estimation of uncertainties in material parameters, optical parameters, etc.

Currently, the coating thermo-optic trace is not an actual calculation; it is just a flat line culled from figure 5.9 of Tara's thesis.

The PDH shot noise trace is shown assuming an incident power of 1 mW on each cavity, a PDH modulation index of 0.2 rad, and a cavity visibility of 0.92.

To do:

  • Finish implementing true TO calculation
  • Add photothermal (requires RIN data)
  • Add seismic (requires seismic data, seismic stack TF data)

I've implemented the TO calculation following Evans et al. (2008), along with the so-called Yamamoto correction for the CTE.

These changes are on the SVN.

Attachment 1: noiseBudget.pdf
noiseBudget.pdf
  1484   Mon Aug 25 03:56:17 2014 taraHowToNoiseBudgetoptimization for ETM with a-Si/SiO2 coatings

 I used optimization codes for ETM. The optimization reduce the PSD of Brownian noise by ~ 3/4 (in units of [m^2/Hz]) from QWL structure.

 Since we have not had all the material parameters for aSi:H at 120K with 1550nm, the optimization here is for room temperature with 1550 nm (for Brownian noise only). 

 opt2_aSi.png

opt2_RT.png

fig1: optical thickness for ETM with minimized BR noise. The transmission is 5.4 ppm and the reflected phase is ~ 179 degree.

Parameters/configuration used in the optimization:

  • T = 300 K   (room temp)
  • wavelength = 1550 nm;
  • Si substrate, n = 3.5;
  • Low index material : fused silica, loss = 0.4e-4, n = 1.444;
  • High index material: aSi:H, loss = 1e-6, n = 3.48; 
  • The coating has SiO2 cap (air-coating surface) for protection
  • Spot radius = 6 cm.
  •  This optimization is only for Brownian noise, we can do another optimization once the thermo-optical properties are known (thermal expansion, dn/dT)

It is remarkable that 5ppm transmission can be achieved with just 17 layers of coatings due to the largely different values between nL and nH. This makes the total thickness down to ~ 3 um.

BR noise from the optimized coating is  3.3x 10^-42 [m^2/Hz] at 100 Hz. This is converted to the strain of ~ 5x10^-25 [1/sqrt Hz] for 4 km interferometer. 

Note: for QWL structure, with 14 layers + half wave cap of SiO2 (total of 15 layers), the transmission is ~5.2 ppm and the coating Brownian noise is 4.2x10^-42 [m^2 /Hz]. So the optimization reduced the PSD of BR noise by ~ 25%. 

  1483   Sun Aug 24 20:07:57 2014 taraNotesVacuumion pump is on

I turned the ion pump for vacuum chamber on. The initial current is 7.3mA ( the value before opening the chamber was 7 uA)
The turbo pump was turned off.

  1482   Wed Aug 20 19:00:03 2014 EvanNotesNoiseBudgetAlGaAs python noise budget

Quote:

I have started a python implementation of the AlGaAs noise budget. All parameters, functions, etc. are defined in a single notebook, and this same notebook generates the plot. The python uncertainties package facilitates estimation of uncertainties in material parameters, optical parameters, etc.

Currently, the coating thermo-optic trace is not an actual calculation; it is just a flat line culled from figure 5.9 of Tara's thesis.

The PDH shot noise trace is shown assuming an incident power of 1 mW on each cavity, a PDH modulation index of 0.2 rad, and a cavity visibility of 0.92.

To do:

  • Finish implementing true TO calculation
  • Add photothermal (requires RIN data)
  • Add seismic (requires seismic data, seismic stack TF data)
Attachment 1: noiseBudget.pdf
noiseBudget.pdf
  1481   Wed Aug 20 12:38:06 2014 Tara, EvanDailyProgressVacuumChamber pumping down

We put on the CF gasket and closed the transmission side of the chamber. Now we are pumping down.

Tara did some work last night to ensure that the window reflections on the input side of the chamber are not overlapping with the cavity reflections. The south window reflection appears to be clipping on the bottom periscope mirror, but we can fix this later.

Next steps:

  • Mode matching (including adjustment of the input lenses)
  • Locking
  • Realignment of transmission optics
  • Re-establishing beat
  1480   Tue Aug 19 10:35:07 2014 EvanDailyProgressRefCavSouth cavity finesse

I swept the south NPRO PZT with a 2 Vpp, 5 Hz triangle wave and watched the transmission of the south cavity using a PDA100A. I saved three such transmission sweeps from the oscilloscope, and then performed Lorentzian fits on each of them in order to get the cavity pole.

From the Lorentzian fits along with the 4.4(2) MHz/V calibration found earlier [and FSR = 4070(30) MHz], I find the following:

  • Lower resonance: cavity pole of 135(3) kHz, corresponding to a finesse of 15100(340). This gives the total round-trip loss as 417(10) ppm.
  • Upper resonance: cavity pole of 139(9) kHz, corresponding to a finesse of 14600(1000). This gives the total round-trip loss as 429(28) ppm.
Attachment 1: southsweep.pdf
southsweep.pdf
Attachment 2: southsweep.zip
  1479   Tue Aug 19 09:25:19 2014 EvanDailyProgressRefCavPolarization selectivity of south AlGaAs cavity

I temporarily removed the QWP immediately before the periscope. Then I added a HWP directly in front of the vacuum chamber window.

While sweeping the laser across the south TEM00 resonances, I monitored the peak voltage of each resonance.

For this particular HWP mount, rotating the mount to 23(1) degrees produces s polarization, in the sense that placing this HWP between two PBSs causes the second PBS to reflect 100% of the beam.

Attachment 1: pol.pdf
pol.pdf
Attachment 2: southpol.zip
  1478   Tue Aug 19 08:55:24 2014 EvanDailyProgressLaserSouth NPRO PZT acutation coefficient

I used a function generator to drive the south NPRO PZT with a triangle wave. Then with the 14.75 MHz sidebands on, I used a PDA100A to watch the south cavity transmission.

Looking by eye at the carrier and sideband transmission peaks, I find an actuation coefficient of 4.4(2) MHz/V, which is higher than what Tara measured in 2010 (maybe the coefficient depends on which axial mode the NPRO is operating on?)

From the attached plot, we can also see that the mode splitting for the south cavity is 2.0(4) MHz.

Attachment 1: south_npro.pdf
south_npro.pdf
Attachment 2: southnpro.zip
  1477   Tue Aug 19 03:55:36 2014 Tara, EvanDailyProgressRefCavSouth cavity OK so far

Quote:

Quote:

Tara has successfully formed the AlGaAs cavities. The configurations are as follows:

  • Spacer 95: to the left of the ATF logo is mirror 114, and to the right of the ATF logo is mirror 143.
  • Spacer 96: to the left of the ATF logo is mirror 141, and to the right of the ATF logo is mirror 132.

Mirror 137 has not been contacted.

Redid optical contacting on south (for a second time) to try to get rid of scattering defects.

Spacer 95: left of ATF logo is 143, right is 137. 143 is on transmission side of chamber.

We redid the mode-matching into south, and judging from CCD images it appears to be free of gross scattering effects.

In the process of moving the seismic stack around, we found that the two rubber noodles on the transmission side had fallen over (so they were being compressed transversely instead of longitudinally). We stood them upright again, but one of them broke, so we had to swap it with a spare. (We tried for a while to make a new one by coring out a cylinder, but they seem to be very brittle. Tara suspects that they're old and broken down.)

Next step is to adjust the stack as necessary to avoid reflections from the windows.

At some point I would like to do the following:

  • Birefringence measurement: temporarily swap QWP before periscope with HWP, record swept transmission as a function of HWP angle
  • Redo NPRO PZT calibration: record swept transmission with 14.75 MHz sidebands on, and thereby infer voltage-to-frequency coefficient
  1476   Mon Aug 18 17:58:57 2014 Tara, EvanDailyProgressRefCavOptical contacting

Quote:

Tara has successfully formed the AlGaAs cavities. The configurations are as follows:

  • Spacer 95: to the left of the ATF logo is mirror 114, and to the right of the ATF logo is mirror 143.
  • Spacer 96: to the left of the ATF logo is mirror 141, and to the right of the ATF logo is mirror 132.

Mirror 137 has not been contacted.

Redid optical contacting on south (for a second time) to try to get rid of scattering defects.

Spacer 95: left of ATF logo is 143, right is 137. 143 is on transmission side of chamber.

  1475   Sat Aug 16 13:06:35 2014 EvanNotesRefCavExpected AlGaAs cavity parameters

Just so we have a concise table that we can refer to:

  North South Note
Length 1.45(1)" = 3.68(3) cm  
FSR 4070(28) MHz c/2L
Mirror ROCs 1.000(3) m Uncertainty is a guess
g factor 0.9632(3) 1 − L/R
TMS 353(3) MHz νFSR×arccos(g)/π
Transmission loss 297(6) ppm 317(3) ppm ctn:1468
Scatter loss 30 ppm 30 ppm Crude guess
Absorption loss 30(30) ppm 30(30) ppm Extremely crude guess
Finesse 16700(1400) 15100(340), 14600(1000) 2π/(losses); ctn:1480
Cavity pole
116(10) kHz 135(3) kHz, 139(10) kHz νFSR/(2F)

 

  1474   Fri Aug 15 15:01:37 2014 Tara, EvanDailyProgressRefCavInserting AlGaAs cavities

Executive summary

  1. We replaced the northeast air spring on the vacuum chamber, because it was leaky.
  2. We opened the transmission side of the vacuum chamber, removed the silica/tantala cavities, and inserted the AlGaAs cavities. The configuration is as follows:
    • SN 00095: south. Logo readable when standing on north side of table.
    • SN 00096: north. Logo readable when standing on north side of table.
  3. We scanned the modes of the north cavity and did some rough mode-matching to TEM00. All modes (including TEM00) appear to be doubled. Is this birefringence?
  4. We scanned the modes of the south cavity. We we able to match into TEM(10)0, then TEM90, TEM80, etc., with relative ease (albeit with the same doubling as observed in the north cavity). However, as we got closer to TEM00, we noticed the presence of two bright scattering centers near the mode axis. These scattering centers appear to be hosing the buildup of the TEM00 mode in the south cavity.
  5. Tara thinks we cannot proceed with the south cavity as is. We'll have to take off and reclean at least one of the mirrors.

Details

At various times, we put the transmission of the north cavity on a PDA100A and monitored the voltage on a scope while sweeping the laser PZT. For the two TEM00 modes of the north cavity, the observed splitting was 11.5 ms when the PZT was driven with a 4 Vpp, 5 Hz triangle wave. Tara has previously measured the south laser PZT actuation coefficient as 3.1 MHz/V (ctn:182). This gives the frequency of the splitting as 1.4 MHz. Since the expected FSR of these cavities is 4070 MHz, this corresponds to a cavity length difference of 180 pm.

The FWHMs of the two peaks (again as seen on the scope) were 1.16 ms and 1.30 ms. With the FSR given above, this gives the finesses as 29 000 and 25 000. That's higher than what should be possible given the measured transmissivities of the mirrors [we expect a finesse 2π/(300 ppm) = 21 000], but this was a quick and dirty measurement that relies on a PZT calibration that's a few years old.

  1473   Thu Aug 14 15:27:36 2014 EvanDailyProgressPMCPMC encap measurements

OD: 1.63"

Depth: ca. 0.9"

Minimum clearance between cap and mount: ca. 0.5"

  1472   Thu Aug 14 15:24:14 2014 Emily NotesopticTemporarily changed angle on half wave plate

(Laser going to ACAV) I changed the angle of the half-waveplate before the PBS in order to increase the amount of power going into the fiber that goes to gyro lab.  Its original position was at 277 degrees.  I put a beam dump behind the lens (PLCX-25.4-38.6-UV-1064) so the higher power does not reach the photodiode.  The new position is at 248 degrees.  I will move it back before I leave.

  1471   Thu Aug 14 15:23:36 2014 Emily, EvanNotesopticAOM fiber noise cancellation

 New setup for fiber phase noise cancellation with one AOM

 
We re-did mode-matching calculations and replaced the lenses before the fiber input in order to optimize the amount of power that comes out of the fiber.  The waist coming out of the PMC is 370 microns.  Following the PMC are the following lenses: placed 7 inches away is a PLCX-25.4.128.8-UV-1064 with a focal length of 250mm, placed 29 inches away is a PLCX-25.4-64.4-C-1064 with a focal length of 125 mm, and placed 35 inches away is a KBX052 with a focal length of 50.2mm. This yields a waist of 69 microns going into the fiber.  Going into the fiber is about 1.1 mw and coming out is approximately 500 micro watts.  We replaced the VCO driver since it was not driving the AOM and had a deformed signal.  Now we are using a Marconi and low-noise amplifier to drive the AOM.  We also replaced the AOM with an Isomet AOM 1205c-843.  
 
We re-did mode-matching calculations into the AOM and to the mirror.  After the fiber output is a waist of 50 microns.  Placed 2 inches away is a: PLCX-25.4-33.7-UV-1064 with a focal length of 50mm, placed 10 inches away is a: PLCX-25.4-77.3-UV-1064 with a focal length of 150mm and placed 18 inches away is a: PLCX-25.4-36.1-UV-1064 with a focal length of 70mm.  The first two lenses before the AOM yield a was it of 150 microns going through the AOM (recommended waist from the Isomet AOM 1205c-843 manual) and the third lens yields a waist of 156 microns at the mirror.  We used a beam dump to block the zeroth order beam, so the only the first order beam is double passed through the fiber.  
 
We are using the same setup to beat the double passed beam with the original beam onto a new focus 1811 photodiode.  The original beam has a power of 850 micro-watts and the double-passed beam has a power of 10 micro-watts. While the efficiency can be improved, for now we will work with what we have in order to prove that our new setup with 1 AOM will cancel the noise in the system.  
 Final_AOM_Setup.pdf
In this setup, we lock the optical beat to the marconi in a PLL.  
The AC signal optical beat fluctuation was 198-428mV.
Once the optical beat was locked to the marconi, we measured the error signal and control signal.  We also measured the control signal without cancellation to make sure that this works.  In order to do the measurement without cancellation, we locked the marconi to the optical beat.  We also measured the open loop transfer function with and without cancellation.  The following data was obtained: 
 
 Unknown-4
Unknown-5
 
 
 
  1470   Sun Aug 10 18:42:48 2014 EvanNotesPMCPMC heater, v2

I'm building this instead:

pmc_heater_2.jpg

  1469   Wed Aug 6 17:29:30 2014 EvanDailyProgressRefCavOptical contacting

Tara has successfully formed the AlGaAs cavities. The configurations are as follows:

  • Spacer 95: to the left of the ATF logo is mirror 114, and to the right of the ATF logo is mirror 143.
  • Spacer 96: to the left of the ATF logo is mirror 141, and to the right of the ATF logo is mirror 132.

Mirror 137 has not been contacted.

Attachment 1: 114.jpg
114.jpg
  1468   Tue Aug 5 18:10:31 2014 EvanDailyProgressopticAlGaAs mirror transmissions; optical contacting

I used the ThorLabs power meter to get the transmission coefficients for the five AlGaAs mirrors.

For each measurement, I wrote down the incident power (20 mW nominal), the transmitted power (≈3.5 µW, depending on the mirror and background light level), and the transmitted power with the beam blocked (to get the dark power).

Mirror
Transmission (ppm) Average (ppm)
#114 142(6) 142(6)
#132 162.4(1.4), 159.8(2.1), 163.0(2.1) 161.7(1.9)
#137 149.8(3.4), 149.5(2.0), 148.0(2.0) 149.1(2.5)
#141

154.9(2.0), 155.4(2.1), 155.4(2.1)

155.2(2.1)
#143 155.6(2.1), 154.7(2.1) 155.2(2.1)

In other news, Tara bonded mirror #114 to spacer #95. The contacting seems to be tough going because of some recalcitrant smudges on the substrate surfaces.

Attachment 1: almost.jpg
almost.jpg
Attachment 2: done.jpg
done.jpg
  1467   Tue Aug 5 14:54:36 2014 EvanDailyProgressopticBRDF of AlGaAs mirror 132 after cleaning

Quote:

Incident power: 20.0(1) mW

Exposure times used: 25 ms, 50 ms, 200 ms, 500 ms, 1000 ms

Transmitted power: 3.34(2) µW. This gives a transmission of 167(1) ppm for this mirror.

TIS from 16° to 73° is 18(1) ppm.

Data and code are on the SVN at CTNLab/measurements/2014_08_05.

 Basically the same story with 132.

Attachment 1: 132brdf.pdf
132brdf.pdf
  1466   Tue Aug 5 08:14:14 2014 EvanDailyProgressopticBRDF of AlGaAs mirror 141 after cleaning

Incident power: 20.0(1) mW

Exposure times used: 25 ms, 50 ms, 200 ms, 500 ms, 1000 ms

Transmitted power: 3.34(2) µW. This gives a transmission of 167(1) ppm for this mirror.

TIS from 16° to 73° is 18(1) ppm.

Data and code are on the SVN at CTNLab/measurements/2014_08_05.

Attachment 1: 141brdf.pdf
141brdf.pdf
  1465   Mon Aug 4 15:23:06 2014 EvanDailyProgressopticBRDF of AlGaAs mirror 114 after cleaning

[Tara, Evan]

Tara also took a BRDF measurement of #114 after cleaning it.

After cleaning, TIS from 14° to 71° is 2.7(5) ppm. Much improved.

Data and code are on the SVN at CTNlab/measurements/2014_07_31.

Attachment 1: 114_cleaned_brdf.pdf
114_cleaned_brdf.pdf
  1464   Sat Aug 2 23:07:39 2014 EvanNotesNoiseBudgetAlGaAs python noise budget

I have started a python implementation of the AlGaAs noise budget. All parameters, functions, etc. are defined in a single notebook, and this same notebook generates the plot. The python uncertainties package facilitates estimation of uncertainties in material parameters, optical parameters, etc.

Currently, the coating thermo-optic trace is not an actual calculation; it is just a flat line culled from figure 5.9 of Tara's thesis.

The PDH shot noise trace is shown assuming an incident power of 1 mW on each cavity, a PDH modulation index of 0.2 rad, and a cavity visibility of 0.92.

Attachment 1: noiseBudget.pdf
noiseBudget.pdf
  1463   Thu Jul 31 09:40:24 2014 EvanDailyProgressopticBRDF of AlGaAs mirror 114

[Tara, Evan]

Tara took a BRDF measurement yesterday of AlGaAs mirror #114.

In this measurement, the return beam is dumped using black anodized foil instead of a razor blade dump. This seems to make the peak at 20° disappear, and now we get a more or less monotonic falloff in scattered power.

TIS from 14° to 71° is 39(6) ppm.

Data and code are on the SVN at CTNlab/measurements/2014_07_30.

Attachment 1: 114brdf.pdf
114brdf.pdf
  1462   Wed Jul 30 17:47:56 2014 taraDailyProgressopticBRDF of AlGaAs mirror 143

 I used the setup to measure scattered loss from an REO mirror (mirror for iLIGO refcav, the one we measured coating thermal noise) and get 6 ppm. This number agrees quite well with the previous Finesse measurement.

 

  Finesse measurement from REO mirrors = 9700 , see PSL:424 The absorption loss in each mirror is ~ 5 ppm ( from photo thermal measurement, see PSL:1375). The measured finesse infers that the roundtrip loss is ~ 24 ppm, see here. So each mirror has ~ 12 ppm loss. With ~ 5ppm absorption loss, we can expect ~ 6-7 ppm loss for scattered loss.  So this measurement roughly says that our scattered light setup and calibration is ok.

 

  1461   Tue Jul 29 11:40:09 2014 EvanDailyProgressopticBRDF of AlGaAs mirror 143

[Tara, Evan]

Yesterday we took a scatter measurement of AlGaAs mirror #143. Instead of one bright scattering center, we saw 3.

The procedure is identical to the procedure used for mirror #137, although the exposure settings and choice of angles are a bit different (see the attached plot). Also, we used 20 mW of incident power instead of 10 mW.

Total integrated scatter from 14° to 82° is 80(8) ppm.

Data, images, and plot-generating code are on the SVN at CTNlab/measurements/2014_07_28.

Attachment 1: 143brdf.pdf
143brdf.pdf
  1460   Sun Jul 27 19:46:16 2014 EvanDailyProgressopticBRDF of AlGaAs mirror 137B1

[Tara, Evan]

We replaced the Lambertian diffuser with AlGaAs mirror 137B1. We intentionally induced a nonzero AOI of the incident beam, so that the reflected beam could be dumped cleanly. At a distance of 25.7(3) cm back from the mirror, the reflected and incident beams were separated by 1.3(1) cm, giving an AOI of 1.45(11)°.

  1. We measured the incident laser power as 9.94(2) mW.
  2. We set the exposure time of the camera to 250 ms.
  3. We swung the boom to 13°, 16°, 19°, 22°, 25°, 28°, 31°, and 34°. At each angle, we took 5 CCD images with the beam incident, and 1 CCD image with the beam blocked.
  4. We measured the incident laser power as 9.95(2) mW.
  5. Because the scattered power had fallen off sharply by 30°, we turned up the exposure time to 1.00 s.
  6. We swung the boom to 31°, 34°, 37°, 40°, and 43°. At each angle, we took 5 CCD images with the beam incident, and 1 CCD image with the beam blocked.
  7. We measured the incident laser power as 10.08(2) mW.
  8. We swung the boom to 46°, 49°, 52°, 55°, 58°, 61°, 64°, 67°, and 70°. At each angle, we took 5 CCD images with the beam incident, and 1 CCD image with the beam blocked.
  9. We measured the incident laser power as 10.06(2) mW.

For all of these measurements, the two ND filters (OD1.5+OD3.0) were not attached; just the RG1000. With the ThorLabs power meter, we measured the combined transmissivity of these two ND filters to be 1865(14) ppm.

The first attachment shows an example CCD image. The second attachment shows the raw counts, the inferred scattered power, and the BRDF.

Attachment 1: ccdImage.pdf
ccdImage.pdf
Attachment 2: 137brdf.pdf
137brdf.pdf
  1459   Fri Jul 25 14:21:34 2014 EvanDailyProgressopticCalibration for scattered light measurement

Quote:

 Do we need to improve this before moving on to the AlGaAs BRDF measurement?

Yes.

We added an OD1.5, an OD3.0, and an RG1000 in front of the camera lens (note that these ODs are probably specked for something other than 1064 nm). Then we increased the exposure time to 20 ms. For the AlGaAs measurement, we may need to increase it even further in order to get good statistics.

Then we fixed the boom at 25° and varied the power using the upstream HWP + PBS combo.

For each power level, we took a measurement with the power meter, then 10 CCD images, then another measurement with the power meter. From this we are able to extract nominal values and uncertanties for the power level and the counts. The result is attached. The calibration has about a 4% uncertainty.

 

Note (Tara): The power measurement includes the solid angle of 3.375 x10^-3 str ( detector diameter = 0.4 inch, distance from the sample = 15.5 cm)

Attachment 1: cal.pdf
cal.pdf
  1458   Fri Jul 25 08:12:40 2014 EvanDailyProgressopticCalibration for scattered light measurement

Quote:

The first attachment is the BRDF of the diffuser based on the power data. The second is the inferred calibration between total CCD counts (with background counts subtracted) and scattered power. The correlation is not great. We may want to retake this data with the room lights off, and also we may want to take multiple exposures per angle setting (that way we can make some estimate of the uncertainty in the CCD counts).

 I put the boom at 15° and took four sets of five exposures. Then I ran my image processing code again to get an uncertainty in the count values. I get the following:

  • Beam incident, room lights on: 546(31) × 103 cts
  • Beam blocked, room lights on: 417(9) × 103 cts
  • Beam incident, room lights off: 547(34) × 103 cts
  • Beam blocked, room lights off: 410(2) × 103 cts

For each set of five, the nominal value is the mean and the uncertainty is the standard deviation of the total counts within the 200×200 pixel region around the beam. Again the exposure time is 100 µs and there was an RG1000 filter in front of the camera lens.

Using a fractional uncertainty of 31/546 = 0.057 for yesterday's background-subtracted total counts, I reran the calibration code. The new plot is attached. The calibration slope (and its uncertainty) doesn't change much, but we can see that the uncertainties in the total counts are quite large. Do we need to improve this before moving on to the AlGaAs BRDF measurement?

Attachment 1: ccdCal.pdf
ccdCal.pdf
  1457   Thu Jul 24 17:11:18 2014 EvanDailyProgressopticCalibration for scattered light measurement

[Tara, Evan, Josh, et al.]

Today we did some characterization and calibration of the scattered light apparatus.

To start with, we examined an AlGaAs mirror (s/n 173). We found that there was a great deal of diffuse light transmitted through the mirror (as seen in fig. 3 of ctn:1456). On Josh's suggestion, we put down an iris about 5" in front of the mirror. We stopped it down just enough so that both the incident and reflected beams could clear the aperture. This made the diffuse stuff disappear.

Next, we swapped out the AlGaAs mirror for a Lambertian diffuser (the same one used in Magaña-Sandoval et al.). Tara affixed the power meter to the camera boom in such a way that it could be raised or lowered in front of the camera.

We adjusted the incident power on the diffuser to be 3.00(1) mW. We then swung the boom in 5° increments from 10° to 70° from normal incidence. At each angle, we took the following:

  1. Power incident on power meter with beam blocked
  2. Power incident on power meter with beam unblocked
  3. CCD image with beam blocked
  4. CCD image with beam unblocked.

The beam was blocked using a dump located immediately upstream of the steering mirror.

The first attachment is the BRDF of the diffuser based on the power data. The second is the inferred calibration between total CCD counts (with background counts subtracted) and scattered power. The correlation is not great. We may want to retake this data with the room lights off, and also we may want to take multiple exposures per angle setting (that way we can make some estimate of the uncertainty in the CCD counts). The third attachment shows the analyzed CCD region for the 10° images; I've restricted the analysis to a 200×200 pixel region around the diffuser.

The exposure time was 100 µs, and there was a 1 µm long-pass filter (RG1000) affixed to the camera lens.

Data, CCD images, and plot-generating code are on the SVN at CTNLab/measurements/2014_07_24.

Attachment 1: power.pdf
power.pdf
Attachment 2: ccdCal.pdf
ccdCal.pdf
Attachment 3: 10.pdf
10.pdf
  1456   Thu Jul 24 02:00:23 2014 taraDailyProgressopticsetting up scattered light measurement

I rechecked the CCD response vs exposure time and power. The results are linear.

 

After some adjustments (strain relief on the camera's cables, clamping down the camera properly), I made sure that the camera is more stable and repeated the measurement. The CCD response is linear with the incident power on the sample (this is under the assumption that the scattered power is directly proportional to the incident power). 

power.png

Fig1:  CCD response vs incident power. The camera response is linear.

== AlGaAs Samples==

I prepared the sample for measurements. All the samples are quite dirty, especially on the flat sides. So I wiped all of them. I still cannot get rid off some water marks on the annulus of the mirror. It might cause some problems when I optical contact the mirrors. I'll try to clean them later.

photo-1.JPG

fig2: one of the AlGaAs mirrors before cleaning. 

I put one of the samples in the scattered light setup. The transmitted beam has a lot of diffused light behind the mirror. The amount of the diffused light changes with the beam direction. I'm not sure exactly why. I'll try to investigate it more. But the scattered light from the sample is very small. Most of the light is from debris on the surface, not the micro roughness of the sample. The amount of scattered light significantly changes with the beam position on the mirror.

photo-2.JPG

fig3: diffused light behind the mirror. It might come from the reflection inside the substrate because the incident beam is not normal to the surface.

  1455   Wed Jul 23 00:28:14 2014 taraDailyProgressopticsetting up scattered light measurement

I'm checking the linearity of power and exposure on the camera. The ccd counts are quite linear with the exposure setup, but I have to check the power again.

 

==ccd count vs exposure setup==

  The exposure time on the camera can be set to adjust the brightness of the image. Since we might have to adjust it to make sure that the images won't be saturated, it is necessary to check if the ccd count response linearly to the exposure setup or not.

I used a silver mirror as a test sample. The incident power is constant, and the camera position is fixed. Then adjust the exposure from 5k to 30k. I'm not sure if it is in nano second or microsecond unit. [Edit, 20140725: according to page 18 of the manual for the Prosilica GC750, the available exposure options are 30 µs to 60 s, in 1 µs increments. —Evan] But from fig1, the ccd count is quite linearly proportional to the exposure value.

exp_vs_light.png

It turns out that when I try to calibrate a sample, the incident power on the sample has to be more (so the power meter can measure some scattered power) and the camera can be saturated. The exposure value has to be around 1000, and I have not checked the response at this level. I might have to remeasure it.

 

==ccd count vs power== 

 This measurement is similar to the above. But this time the incident power (to the sample) is varied. The result is not linear. I check the images and see that the bright spot moves. The camera might move during the measurement. I'll repeat this again. It will be complicated for the calibration if the ccd count is not linear with the power.

power.png

== To do==

  • check the ccd count for exposure value down to lowest setting.
  • check the ccd count for different power incident.
  • check the ccd count with different ND filter in front of the camera.

 

  1454   Tue Jul 22 23:55:08 2014 Emily, EvanNotesNoiseBudgetPSL VCO noise

VCO driver noise measurement

To measure the noise of the PSL VCO driver, we used the same PLL set-up from previous noise measurements.  The PLL consisted of the following: IFR/Marconi 2023A, the SR560, Mini-Circuits Frequency Mixer ZX05-1MHW-S+-0.5-600MHz, Mini-Circuits 15542 BLP-5+ Low Pass Filter 50 Ohm DC-5MHz, and the Stanford Research Systems Model SR560- Low Noise Pre-amplifier with a gain of 200 V/V.  We connected another VCO to the RF port of the mixer.   The Marconi had a carrier frequency of 80 MHz, an RF level of 13 dBm and FM dev set to 1 KHz Ext. DC.  

We connected the VCO to a power supply by hooking up a 9-pin dsub breakout box into the VME interface.  The VCO driver needs 24V from the power supply.  From opening up the box, we found that there are three test points in the VME interface: TP1, TP2 and TP3.  TP1 corresponds to -24V, TP2 corresponds to +24V and TP3 is ground.  Additionally, we needed to figure out what pins to hook up the positive, negative and ground cables onto the breakout box.  +24 V corresponds to pins 9 and 4, -24 V corresponds to pin 5 and ground corresponds to pins 8 and 3.  There are also two switches that need to be connected to the ground in order for the driver to function properly.  The test switch, which corresponds to pin 1 and the wide switch, which corresponds to pin 6 are both connected to pin 3(ground).  We used the TENMA Laboratory DC Power Supply 72-2080 and it was set to 24 V and .5 A.  

 

After locking the frequencies, we measured the transfer function and ASD with an FFT analyzer (Agilent 35670A Dynamic Signal Analyzer).  The following data was obtained:

 PLL_VCO.epsBudget_VCO.png

  1453   Fri Jul 18 12:51:19 2014 Emily, EvanNotesNoiseBudgetMarconi noise

Marconi Noise Measurement

 To verify whether or not the noise was from the AOM driver, we measured the noise in the Marconi.   We set up a PLL to do this measurement.  We used the IFR/Marconi 2023A, the SR560, Mini-Circuits Frequency Mixer ZX05-1MHW-S+-0.5-600MHz, Mini-Circuits 15542 BLP-5+ Low Pass Filter 50 Ohm DC-5MHz, and the Stanford Research Systems Model SR560- Low Noise Pre-amplifier with a gain of 500 V/V.  We  connected another Marconi to the RF port of the mixer.   Marconi 1 had a carrier frequency of 80 MHz, an RF level of 13 dBm and FM dev set to 1 KHz Ext. DC.  The second Marconi that we used had a carrier frequency of 80 MHz, an RF level of 2 dBm to match that of the AOM driver and FM dev disabled to prevent noise.  

 Set-up:

 PLL3.eps

Measurement: After locking the frequencies, we looked at the measured the PSD on an FFT analyzer (Agilent 35670A Dynamic Signal Analyzer) and obtained a measurement of the ASD. Unknown.png

It appears that the noise is coming from the AOM driver.  

 

  1452   Thu Jul 17 18:57:54 2014 taraDailyProgressopticsetting up scattered light measurement

I'm testing the setup and a code for extracting scattered light from the images. 

 

I used a red laser pointer to test the scattered light setup. Then took a picture with no light (fig1) and a picture with the incident light (fig2). The scattered light can be extracted by subtract fig1(background) from fig2.

The snapshots saved by SampleViewer are in .bmp file. When it is read by MATLAB, the file will contain 480x752x3 matrix element, Each are varied between 0 and 255. The values are proportional to the brightness (how many photons hit the cell). 480x752 is the resolution of the image, x3 are for R G B color. In our case, the image is greyscale and the values are identical. The code can be found in the attached file.  

 dark1.png

fig1: The test mirror without incident beam taken as a background image. The image is enhanced by a factor of 5 (by matlab).

 

light.png

fig2: The test mirror with a red incident beam around the center. The image is enhanced by a factor of 5.

 

residual.png

fig3: the image is created by subtracting data of fig1 (background) from fig2 (scattered light) and enhanced by a factor of 100. The scattered light on both surfaces can be seen clearly around the center.

 ==To do next==

  • From fig 3, the background can be seen even after subtraction, so some black curtains and beam dumps should be added behind the mirror.
  •  A room light filter should be installed in front of the camera.
  • I'll see if we can find a sample with known scattering loss, so that we can compare how accurate the measurement is.
Attachment 4: scattered_.zip
ELOG V3.1.3-